OpenStudy (anonymous):

write an equation for each sequence below and determine the 8th term. d) 7, 7.5, 7.75, 7.875...

OpenStudy (anonymous):

7+ (0.5)^x

OpenStudy (anonymous):

x begins from 0

OpenStudy (anonymous):

x being the nth term? but that doesn't work... if x starts from 0 then that me ants the first term should be 8 not 7.

OpenStudy (anonymous):

when x=0 then ans will be 8

OpenStudy (anonymous):

see modify it to 7 + (0.5)^(n-1)

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

it still makes the first term 8 not 7.

OpenStudy (anonymous):

8th term will be 7.9921875

OpenStudy (anonymous):

trying to make an equation

OpenStudy (rulnick):

nth term is 8 - (1/2)^(n-1), and 8th term is 8 - (1/2)^(7) = 8-1/128 = 7 127/128.

OpenStudy (anonymous):

that's what i got as well. this is the equation my teacher came up with when i asked him \[t _{n}=8-1/2^{n-1}\]

OpenStudy (anonymous):

ye i got 7.992

OpenStudy (anonymous):

yes this equation perfectally suit the series

OpenStudy (anonymous):

okay. this is the right equation... i just don't get how he got to that. he told me that \[t _{1}= 7\] \[t_{2}= 7 + 1/2 \] \[t_{3}= 7 + 1/2 + 1/2^{2}\] and so on. and then he got a =1/2, n=n-1, r=1/2 we know that a geometric sequence equation is represented by \[t_{n} = ar^{n-1}\] therefore with that, he did \[t_{n}= 7 + 1/2(1-1/2^{n-1})/1-1/2\] \[=8-1/2^{n-1}\]